# Learning from Emotions, Demographic Information and Implicit User Feedback in Task-Oriented Document-Grounded Dialogues

### v1

@@ Abstract
The success of task-oriented and document-grounded dialogue systems depends on users accepting and enjoying using them. To achieve this, recently published work in the field of Human-Computer Interaction suggests that the combination of considering demographic information, user emotions and learning from the implicit feedback in their utterances, is particularly important. However, these findings have not yet been transferred to the field of Natural Language Processing, where these data are primarily studied separately. Accordingly, no sufficiently annotated dataset is available. To address this gap, we introduce FEDI, the first English dialogue dataset for task-oriented document-grounded dialogues annotated with demographic information, user emotions and implicit feedback. Our experiments with FLAN-T5, GPT-2 and LLaMA-2 show that these data have the potential to improve task completion and the factual consistency of the generated responses and user acceptance111 Code and data is available in GitHub..
Dataset Source Type Demographics Emotions
Implicit
@@
Table 1: Comparison of FEDI to other datasets that provide related annotations. FEDI is comparable with other LLM-generated datasets in terms of avg. turn and utterance length, and has a higher lexical diversity (based on MTLD McCarthy and Jarvis (2010) [^3] We used the Python package lexical-diversity v0.1.1 for calculation (last accessed 04 January 2024).) than many of the crowdsourced datasets.
@@ 1 Introduction
Figure 1: A feedback dialogue from FEDI. User emotion and implicit feedback annotations (generation error and user reaction types) are beneath the utterances.
Task-oriented and document-grounded dialogue systems, such as personal assistants or customer service agents, often implement sensitive tasks, which is why user acceptance and trust are important success criteria. Recently published work in the field of Human-Computer Interaction (HCI) suggests that considering demographic information (such as age, occupation or language style), interpreting user emotions, and learning from implicit user feedback, such as a correction or question in response to a generation error in the preceding system utterance, are particularly important because they often interact and correlate with each other Pelau et al. (2021); Araujo (2018); Zamora (2017); Ciechanowski et al. (2019). For example, in the second utterance of Figure 1 the system misinterprets Claudia’s question, resulting in an incorrect response. This has a direct impact on her emotional state and she now asks the system for clarification (implicit feedback). After the system has generated a more suitable response, she is satisfied, her emotional state changes again and she asks the system another question.
As of today, these findings have not yet been transferred to research in the field of Natural Language Processing (NLP), where these data are usually considered separately from each other. As a result, a large number of works are available that either focus on learning from demographic information Lee et al. (2022); Zhang et al. (2018), user emotions Firdaus et al. (2020); Rashkin et al. (2019); Hsu et al. (2018); Hwang et al. (2023) or implicit user feedback Hancock et al. (2019); Veron et al. (2021); Xu et al. (2023a, b); Ung et al. (2022), but none that considers them together, which is a limitation for future research Chaves and Gerosa (2021); Minjin Rheu and Huh-Yoo (2021).
In this work, we address this gap by (a) providing FEDI, the first English dialogue dataset annotated with implicit user Feedback, Emotions and Demographic Information, and (b) investigating the effect of including these data as part of the input on the performance of three different state-of-the-art language generation models, i.e., FLAN-T5 Chung et al. (2022), GPT-2 Radford et al. (2019) and LLaMA-2 Touvron et al. (2023). FEDI consists of overall 8k task-oriented document-grounded dialogues, 6k of which are annotated with implicit user feedback.
To create FEDI in a cost and resource efficient way, we use GPT-3.5-Turbo [^4] OpenAI GPT-3.5 Model Page (last accessed on 02 January 2024). The model is based on Ouyang et al. (2022). The data was generated between March and June 2023. to generate and annotate the training and validation data, and recruit humans to assess its quality and to collect a separate set of test dialogues, due to the potential limitations of synthetic data Ji et al. (2023); Zhang et al. (2023); Malaviya et al. (2023). We also use GPT-3.5-Turbo for the generation of implicit user feedback. In contrast to previous works which usually focus on feedback specific to their use case Xu et al. (2023b); Ung et al. (2022); Hancock et al. (2019), we use the taxonomies proposed by Petrak et al. (2023) for this, since they cover a variety of different feedback situations, such as feedback on ignored user requests, knowledge gaps or harmful social behavior. In summary, we provide these contributions:
@@
    FEDI, the first task-oriented document-grounded dialogue dataset for learning from demographic information, user emotions and implicit user feedback.
@@
    New experimental insights showing that learning from demographic information, user emotions and implicit feedback has a positive impact on task completion and factual consistency of generated responses in task-oriented knowledge-grounded dialogues.
@@
    A new framework for generating and annotating high-quality task-oriented knowledge-grounded feedback-annotated dialogue data.
FEDI is comparable to other related human-bot and LLM-generated datasets in terms of size, lexical diversity and dialogue length (see Table 3). In our analysis, we provide insights into the quality of generated annotations, especially in comparison to the test data collected by humans.
@@ 2 Related Work
Demographic information, user emotions and implicit user feedback are usually considered separately in the field of NLP. However, available works show that considering demographic information on dialogue-level and as part of the dialogue context has a positive impact on the user experience Lee et al. (2022); Zhang et al. (2018); Hwang et al. (2023); Siddique et al. (2022); Luo et al. (2019). This is similar for user emotions, which are usually considered on utterance-level and detected by separate modules, such as emotion detectors Firdaus et al. (2020); Rashkin et al. (2019); Hsu et al. (2018). To improve upon implicit user feedback usually requires to train the model with the feedback data Ung et al. (2022); Xu et al. (2023a); Veron et al. (2021). In this regard, continual learning has shown to be very promising Xu et al. (2023b); Wang et al. (2019); Hancock et al. (2019). Table 3 gives a comparison of the datasets resulting from these works. It shows that (1) no dataset is available for task-oriented dialogues, and (2) most of the available datasets are the result of resource-intensive crowdsourcing efforts, often leading to datasets of varying quality due to, e.g., methodological artifacts or annotator biases Yang et al. (2023); Parmar et al. (2023); Thorn Jakobsen et al. (2022); Prabhakaran et al. (2021). As an alternative, recent works suggest synthetic data generation using large language models as a more efficient approach to generate high-quality dialogue data Kim et al. (2023); Li et al. (2023); Lee et al. (2022), despite their limitations, such as a tendency to generate hallucinated or harmful output Ji et al. (2023); Zhang et al. (2023); Malaviya et al. (2023).
In this work, we use GPT-3.5-Turbo to generate a task-oriented document-grounded dialogue dataset annotated with demographic information, user emotions and implicit feedback to investigate the impact of such information on the performance of three state-of-the-art language generation models. Due to the potential limitations of synthetic data, we recruit human annotators for quality assessment, curation, and to collect a separate set of test dialogs.
Figure 2: Overview of our framework for generating and annotating dialogues. We distinguish feedback-free and feedback dialogues. Feedback scenarios and error and user reactions types (marked with a *) are specific to feedback dialogues. We provide GPT-3.5-Turbo with all the information from the box on the left side to (1) generate the dialogue in Step 1 and (2) generate the slot and emotion annotations in Step 2.
@@ 3 FEDI
FEDI covers four use cases for task-oriented document-grounded dialogue systems from three domains, including post office services, receptionist services and customer services in the insurance domain. For post office services, we include (1) customer support for parcel shipping, i.e., guiding them through the process of parcel shipping from choosing the right shipping box to informing them about the approximate delivery time, and (2) topping up a prepaid SIM card. For reception and customer services in the insurance domain, we include one use case each, i.e., access control (the reception and registration of new visitors in office buildings) and question answering (in the context of financial topics and pet, health and heritage insurance [^5] We use the policies of POSTE Italiane as a reference for this, which are also publicly available in English (last accessed 04 January 2024).). The question answering dialogues are additionally annotated with the documents that provide the knowledge required for response generation. Appendix A describes the tasks in more detail, including slots, intents and examples.
Demographic Information
We consider gender, age, occupation, name, and language style as demographic information in this work. Overall, we distinguish 12 different language styles, such as formal, dialect and jargon, five demographic cohorts, ranging from Boomers (born between 1952 and 1962) to Generation Alpha (born between 2007 and 2016), and a variety of $1,155$ occupations and $1,000$ names, sampled from The Gazette [^6] Available in GitHub (last accessed on 04 January 2024).. We provide more details in Appendix B.
User Emotions
For user emotions, we use the taxonomy from EmotionLines Hsu et al. (2018), which covers seven different emotions, including Neutral, Joy (which we refer to as Happiness), Sadness, Surprise, Fear, Anger, and Disgust. We extend this list with four emotion types we found to be relevant in related work Kim et al. (2023); Rashkin et al. (2019), including Confusion, Curiosity, Frustration, and Stress. We consider Confusion, Frustration, Fear, Sadness, Disgust, Stress, and Anger as negative emotions.
Implicit User Feedback
For the generation and annotation of implicit user feedback, we use the user reaction type taxonomy proposed by Petrak et al. (2023), which distinguishes five user reaction types in response to generation errors in preceding system utterances, including (1) Ignore and Continue, (2) Repeat or Rephrase, (3) Make Aware With Correction, (4) Make Aware Without Correction, and (5) Ask for Clarification. (2), (3) and (5) describe user reactions that contain implicit user feedback. For generation errors in system utterances, they also propose an error taxonomy of ten types, nine of which are relevant for task-oriented document-grounded dialogues, such as Ignore Request, Attribute Error, Factually Incorrect, or Lack of Sociality. We only consider these error types for generating implicit user feedback. Definitions, further details and examples can be found in Appendix B.
@@
Depending on whether knowledge from a document $D_{t}$ is required to generate $S_{t}$ or the user emotion $E_{t}$, demographic information $DI$, generation error $GE_{t}$, or implicit user feedback $F_{t}$ should be considered, $K=\{D_{t},DI,E_{t},GE_{t},F_{t}\}$. $DI$ includes the user’s gender, age range, occupation, name, and language style. Belief state $B_{t}$ includes the slot values inferred from the dialogue context $C$, which may be used to query knowledge from an external information retrieval system Chen et al. (2022); Hosseini-Asl et al. (2020), such as the document $D_{t}$ in the case of customer service or registration information in the case of access control. We do not use such an additional retrieval system in this work and provide all the information required to generate the response $S_{t}$ as part of the additional information $K$ in the input sequence.
@@ 4 Framework for Generating and Annotating Dialogues
Figure 2 gives an overview of our framework for generating and annotating dialogues using GPT-3.5-Turbo. In general, we distinguish feedback-free and feedback dialogues, i.e., dialogues that provide annotations for implicit user feedback. However, the procedure for dialogue and annotation generation is in general the same for both (Section 4.1). Section 4.2 describes the steps specific to feedback dialogues. For each step that involves GPT-3.5-Turbo, we require the model to return the results in a predefined JSON scheme. If in one step the generation does not match this requirement, does not provide all required fields or contains invalid encoding, the whole dialogue is discarded. We provide more details, including the instructions used in this procedure, in Appendix C.
@@ 4.1 General Approach to Dialogue Generation
For dialogue generation, we provide GPT-3.5-Turbo with the randomly sampled demographic information for the user, a task description, and the role of the starting actor, i.e., user or system. As indicated by the illustration on the left side of Figure 2, a task description describes the flow of events and information which needs to be conveyed by each role to fulfill the task. In the case of question answering, the task description also includes a randomly sampled list of documents from the respective topic. Similar to PersonaChatGen Lee et al. (2022), we instruct the model to use the task description and the demographic information to generate a background story to guide the conversation, such as depicted in the center of Figure 2. We also instruct the model to return the utterance-level annotations for intents and limit the dialogue to 13 turns, since we found that longer dialogues tend to deviate from the task description. For background stories, we limit the length to five sentences, since we found that longer background stories tend to be distracting.
@@
The generation of annotations is the second step in our generation framework [^7] We also tried to reduce API calls by combining dialogue and annotation generation but found that this does not produce reliable results. (see Figure 2). For slot annotations, we provide GPT-3.5-Turbo with the generated dialogue and a list of all slots defined in the task description, possible values and examples. We also instruct the model to only assign and copy values from the dialogue (to prevent hallucinations) and to return the annotations on utterance-level. For emotion annotations, we instruct the model to predict the emotion for each user utterance in the dialogue, given the dialogue and our emotion taxonomy.
@@ 4.2 Feedback Dialogues
For each feedback dialogue, we first use GPT-3.5-Turbo in a prior step to generate three feedback scenarios that are then used as an additional source for dialogue generation. A feedback scenario describes a generation error and the following implicit user feedback (the center of Figure 2 shows an example) [^8] We generate all feedback scenarios for a dialogue at once, using the same API call.. For this, we provide the model with the task description and a list of randomly sampled error and user reaction types. To ensure coherence, we add the constraint that feedback scenarios must be non-mutually exclusive and together form a story in the context of the task description. We limit the number of feedback scenarios per dialogue to not exceed the length of 13 turns per dialogue and to leave room for the start and end of the conversation.
@@
For feedback dialogue generation, we instruct GPT-3.5-Turbo to consider each feedback scenario in three utterances in the generated dialogue, in addition to the constraints described in Section 4.1: The system utterance with the generation error, a subsequent user utterance that reflects the user reaction, and a following system utterance that addresses the user reaction. Next, we generate three additional versions of the same dialogue, each resolving one of the feedback scenarios. Figure 3 illustrates this procedure.
Figure 3: During the generation of feedback dialogues, we subsequently resolve feedback scenarios in the original feedback dialogues.
We first mask the system utterance with generation error and generate a replacement using the task description and the preceding dialogue context. Next, we drop the following two utterances, since they are directly related to the original generation error. This way, the conversation continues with the next regular user utterance. We continue the process until all feedback turns have been resolved as in Version 4.
@@ 5 FEDI Analysis
FEDI consists of 8,852 dialogues, divided into 1,988 feedback-free dialogues, including 326 test dialogues, and 6,864 feedback dialogues (1,716 in four versions, each with one feedback scenario less per dialogue). The test dialogues were collected human-human by eight computer science students in overall 136 paid working hours (see Appendix D for the procedure and details on the hiring process and salary). In the following, we focus on the completeness of generated slot and intent annotations, the distribution of user emotions and the feedback scenarios represented in the dialogues. We provide additional statistical analysis in Appendix E, including split sizes and the distribution of demographic information.
@@
Table 10 shows the ratio of dialogues for which intent and slot annotations were successful, i.e., dialogues that provide all annotations for intent and required slot values.
@@
Dialogues
Feedback Dialogues
Gen. Test Version 1 Version 2 Version 3 Version 4
@@
Table 2: The table shows the ratio of dialogues that are complete in the sense that they are annotated with all intent and slot values [^10] Hallucinated slot values, i.e. slot annotations that do not occur in the respective utterance, are counted as missing..
For the feedback-free dialogues, we distinguish between generated dialogues (Gen.) and test dialogues (Test, collected by humans). We observe large differences between (1) question answering and the other tasks and (2) generated dialogues and the test dialogues collected by humans. We found that this is mostly due to variations in the slot annotations. While the slot annotation scheme for question answering is rather simple (see Appendix A), this is different for the other tasks where slots often depend on the the background story. For example, in the case of parcel shipping, if the user already has a shipping box and just requires information on the shipping procedure, details about available shipping box types are negligible. While human annotators take this into account and occasionally omit slots that are not required based on dialogue motivation, GPT-3.5-Turbo follows our instructions more closely, which include all slots as part of the task description. For feedback dialogues, we observe that the generated corrections not always address the missing information required by the task description. We provide more analysis as part of our human curation study in Section 6.
@@
As expected, negative emotions are more common in feedback dialogues. For Curiosity, we found that the polarity depends on the dialogue context, e.g. whether the previous system utterance successfully addressed the user’s request. It is an emotion that can be either positive or negative, thus it is frequently observed in both dialogues types. Happiness in feedback dialogues is mostly observed as a reaction to system utterances that implement user feedback.
@@
Figure 5 shows the distribution of user reactions in relation to error types represented in the feedback scenarios of the feedback dialogues.
Figure 5: Distribution of user reactions in relation to error types represented in feedback scenarios.
The figure shows that our approach for generating feedback scenarios resulted in meaningful combinations of error and user reaction types. For example, Factually Incorrect is mostly addressed by Make Aware with Correction. Ignore and Continue are only rarely observed. The same applies to Unclear Intention and Attribute Error, which are mostly addressed by Ask for Clarification and Repeat or Rephrase. The latter one is also frequently observed in combination with Ignore Question and Ignore Expectation errors.
Experiment Matching Task Completion Quality
F1 BLEU BertScore Inform Success Intent Acc Slot Acc Toxicity
FLAN-T5 Feedback-Free FLAN-T5 45.0 20.0 88.3 86.7 85.9 54.8 60.9 0.02 52.7
 +Emotions 46.7 (+1.7) 21.0 (+1.0) 88.9 (+0.6) 83.9 (–2.8) 83.2 (–2.7) 61.2 (+6.4) 58.3 (–2.6) 0.02 57.5 (+4.8)
 +Demographics 43.2 (–1.8) 18.4 (–1.6) 87.7 (–0.6) 87.0 (+0.3) 86.0 (+0.1) 33.5 (–21.3) 29.3 (–31.6) 0.03 (+0.01) 54.5 (+1.8)
 +Emotions
 +Demographics
44.2 (–0.8) 19.1 (–0.9) 88.1 (–0.2) 85.3 (–1.4) 85.1 (–0.8) 43.9 (–10.9) 36.7 (–24.2) 0.02 56.4 (+3.7)
Feedback  +Generation Error 41.4 (–3.6) 19.8 (–0.2) 87.8 (–0.5) 96.8 (+10.1) 92.7 (+6.8) 72.5 (+17.7) 76.7 (+15.8) 0.02 56.9 (+4.2)
 +User Reaction 41.3 (–3.7) 19.3 –0.7 87.6 (–0.7) 96.6 (+9.9) 94.1 (+8.2) 69.0 (+14.2) 76.2 +15.3 0.02 56.3 (+4.1)

### v2

# Learning from Implicit User Feedback, Emotions and Demographic Information in Task-Oriented and Document-Grounded Dialogues

@@ Abstract
Implicit user feedback, user emotions and demographic information have shown to be promising sources for improving the accuracy and user engagement of responses generated by dialogue systems. However, the influence of such information on task completion and factual consistency, which are important criteria for task-oriented and document-grounded dialogues, is not yet known. To address this, we introduce FEDI, the first English task-oriented and document-grounded dialogue dataset annotated with this information. Our experiments with Flan-T5, GPT-2 and Llama 2 show a particularly positive impact on task completion and factual consistency. Participants in our human evaluation reported that the responses generated by the feedback-trained models were more informative (Flan-T5 and GPT-2), relevant and factual consistent (Llama 2).111 Code and data are available in GitHub.
[…]
Learning from Implicit User Feedback, Emotions and Demographic Information in Task-Oriented and Document-Grounded Dialogues
Dominic Petrak  and Thy Thy Tran and Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab), Department of Computer Science and Hessian Center for AI (hessian.AI), Technical University of Darmstadt www.ukp.tu-darmstadt.de
Dataset Source Type
Demographic
Information
User
Emotions
Implicit User
@@
Table 1: Comparison of FEDI to other datasets that provide related annotations. FEDI is comparable to other synthetic datasets generated by large language models (LLMs) in terms of avg. turn and utterance length. It also has a higher lexical diversity than many of the crowdsourced datasets [^3] We used the Python package lexical-diversity v0.1.1 for calculation (last accessed 04 January 2024), which implements the approach proposed by McCarthy and Jarvis (2010)..
@@ 1 Introduction
Figure 1: A feedback dialogue from FEDI, annotated with user emotions and implicit user feedback (generation error and user feedback types).
Implicit user feedback Xu et al. (2023b); Veron et al. (2021); Hancock et al. (2019), such as clarification questions, user emotions Hwang et al. (2023); Rashkin et al. (2019); Hsu et al. (2018) and demographic information Lee et al. (2022); Zhang et al. (2018), such as age or language style, are promising sources for improving the accuracy and user engagement of responses generated by dialogue systems. For example, in the second utterance of Figure 1, the system generates a response unrelated to the user’s question, which affects her emotional state. She asks the system for clarification, getting a more satisfying response. This makes her happy and she continues the conversation. However, we do not know to what extent the generated response contributes to achieving the user’s goal and reflects the underlying knowledge source. This is commonly referred to as task completion and factual consistency. Both are important criteria for task-oriented and document-grounded dialogue systems Nekvinda and Dušek (2021); Honovich et al. (2021); Budzianowski et al. (2018), but the impact of implicit user feedback, emotions and demographic information on them is an open research question.
To address this gap, we introduce the FEDI dataset. Following recent research that includes information-seeking in task-oriented dialogues Taranukhin et al. (2024); Braunschweiler et al. (2023); Feng (2021); Campos et al. (2020), e.g., for handling multi-domain scenarios, FEDI provides annotations for required knowledge documents and is the first English task-oriented and document-grounded dialogue dataset annotated with implicit user Feedback, Emotions and Demographic Information. FEDI allows us to investigate the impact of this information on task completion and the factual consistency of responses generated by dialogue systems, for which we use Flan-T5 Chung et al. (2022), GPT-2 Radford et al. (2019) and Llama 2 Touvron et al. (2023b) in this work. We use GPT-3.5444We used GPT-3.5-Turbo (OpenAI GPT-3.5 Model Page, last accessed on 02 January 2024). The model is based on Ouyang et al. (2022). The data was generated between March and June 2023. to generate and annotate the training and validation data for FEDI. We recruit humans to assess its quality and to collect a separate set of test dialogues. In summary, we provide these contributions:
@@
    New experimental insights on the impact of learning from implicit user feedback, user emotions and demographic information, including task completion and factual consistency, and how humans perceive the responses generated by the resulting models.
@@
    FEDI, the first task-oriented and document-grounded dialogue dataset for learning from implicit user feedback, emotions and demographic information. It is comparable to other related datasets in terms of size, lexical diversity and dialogue length (see Table 3).
@@
    A framework for generating and annotating task-oriented and document-grounded feedback-annotated dialogue data. Our analysis provides insights into the quality of the generated dialogues annotations.
@@ 2 Related Work
Learning from user emotions and demographic information can improve generation accuracy and user engagement in dialogue systems Feng et al. (2022); Hsu et al. (2018); Siddique et al. (2022); Zhang et al. (2018). The same applies to implicit user feedback, which usually requires user interaction for data collection and continual learning Xu et al. (2023b); Ung et al. (2022); Wang et al. (2019); Hancock et al. (2019). Table 3 provides a concise overview of the related datasets. None of them contains annotations for all three signals, and most of them were collected in resource-intensive crowdsourcing efforts. This is particularly complex in the context of user feedback Xu et al. (2023b); Ung et al. (2022) and no guarantee for quality Parmar et al. (2023); Yang et al. (2023); Thorn Jakobsen et al. (2022); Prabhakaran et al. (2021). Although LLMs are heavily dependent on detailed instructions and still tend to generate biased, hallucinated, or harmful data Yang et al. (2023); Ji et al. (2023); Zhang et al. (2023b); Malaviya et al. (2023), recent works suggest these models, especially GPT-3.5, as a more efficient approach to generate dialogue data Stricker and Paroubek (2024); Kim et al. (2023); Li et al. (2023); Lee et al. (2022).
To investigate the impact of implicit user feedback, emotions and demographic information in task-oriented and document-grounded dialogues, we create FEDI by combining the best of both worlds. We use GPT-3.5 to generate training and validation data and recruit human annotators for dialogue quality assessment, annotation curation, and collection of a separate set of test dialogues.
@@ 3 FEDI
Self-service terminals are increasingly common in the service sector, including postal services, access controls (e.g., to security-critical areas or in the hospitality industry), or customer service Abbate et al. (2024); Tuomi et al. (2021); Lee et al. (2010). They implement specific workflows to serve customers and support employees. FEDI covers four use cases from these domains. For postal sevices, we include customer support for parcel shipping and topping up a prepaid SIM card. For receptionist and insurance services, we include one use case each, i.e., access control (the reception and registration of new visitors in office buildings) and question answering (in the context of financial topics and pet, health and heritage insurance). The question answering dialogues are additionally annotated with knowledge documents. Appendix A describes the tasks in more detail, including slots (information required for task completion), intents and document sources.
Implicit User Feedback ($GE$, $F$)
We use the taxonomies proposed by Petrak et al. (2023) to generate and annotate generation errors ($GE$) and subsequent implicit user feedback ($F$). They distinguish ten types of generation errors. Nine of which are relevant for FEDI, such as Attribute Error, Factually Incorrect or Lack of Sociality. For implicit user feedback, they distinguish five types, e.g., Ask for Clarification, Ignore and Continue and Repeat or Rephrase. Definitions, further details and examples can be found in Appendix B.
Demographic Information ($DI$)
We consider gender, age, occupation, name, and language style as demographic information in this work. Overall, we distinguish 12 different language styles, such as formal, dialect and jargon, five demographic cohorts, ranging from Boomers (born between 1952 and 1962) to Generation Alpha (born between 2007 and 2016), a variety of $1,155$ occupations, and $2,000$ names. We provide more details, including data sources in Appendix B.
User Emotions ($E$)
Inspired by Hsu et al. (2018), Rashkin et al. (2019) and Kim et al. (2023), we derive a taxonomy of 11 emotions that are potentially relevant for our dialogue tasks, including Neutral, three positive emotions (Curiosity, Surprise and Happiness) and seven negative emotions (Confusion, Frustration, Fear, Sadness, Disgust, Stress, and Anger).
@@
Depending on whether knowledge from a document $D_{t}$ is required to generate $S_{t}$ or the user emotion $E_{t}$, demographic information $DI$, generation error $GE_{t}$, or implicit user feedback $F_{t}$ should be considered, $K=\{D_{t},DI,E_{t},GE_{t},F_{t}\}$. $DI$ includes the gender, age range, occupation, name, and language style of the user. Belief state $B_{t}$ includes the slots predicted for user utterance $U_{t}$.
[…]
@@ 4 Framework for Generating and Annotating Dialogues
Figure 2 gives an overview of our framework for generating and annotating dialogues. We distinguish feedback-free and feedback dialogues, i.e., dialogues that provide annotations for generation errors and implicit user feedback. For each step, we require GPT-3.5 to return the results in a predefined JSON scheme. If in one step the generation does not match this requirement, the whole dialogue is discarded. We provide more details, including the instructions used in this procedure, in Appendix C.
@@ 4.1 General Approach to Dialogue Generation
The procedure is basically the same for feedback-free and feedback dialogues. It starts in the second box from the left in Figure 2. We provide GPT-3.5 with randomly sampled demographic information for the user and a task description, which describes the flow of events to fulfill the task, including the role of the starting actor, i.e., user or system, and a randomly sampled list of documents in the case of question answering. Feedback dialogues require feedback scenarios as additional sources (Section 4.2). We instruct the model to use the task description and demographic information to generate a background story to guide the conversation Lee et al. (2022); Stricker and Paroubek (2024), such as depicted in the center of the figure. We require the model to return the utterance-level annotations for intents (not included in the figure) and limit the dialogue to 13 turns, since we found that longer dialogues tend to deviate from the task description. We limit the length of background stories to five sentences to avoid them becoming a distraction.
@@
For slot annotations, we provide GPT-3.5 with the generated dialogue and the task description (including a list of all slots, possible values, and examples [^5] We also tried to reduce API calls by combining dialogue and annotation generation but found that this does not produce reliable results.). We instruct the model to only copy values from the dialogue and to return the annotations on utterance-level. For emotion annotations, we provide the model with the emotion taxonomy (instead of the task description) and instruct it to predict the emotion for each user utterance.
@@ 4.2 Feedback Dialogues
A feedback scenario describes a generation error and the following implicit user feedback. Figure 2 shows an example in the second box from the left. For generation (first box), we provide GPT-3.5 with the task description and a list of randomly sampled generation error and implicit user feedback types. To ensure coherence, feedback scenarios must not be mutually exclusive and together form a story in the context of the task description. For each feedback dialogue, we generate three feedback scenarios that are then used as an additional source for dialogue generation [^6] We generate all feedback scenarios for a dialogue at once, using a single API call..
@@
For feedback dialogues, we instruct GPT-3.5 to consider each feedback scenario in three consecutive utterances in the generated dialogue: First, the system utterance with the generation error, e.g., Yes, I can help you send a parcel to Paris. Then the subsequent user utterance, e.g., No, the destination is London, not Paris!, which we consider as implicit user feedback. Finally, the following system utterance that addresses the user feedback, e.g., Apologies for the mistake. Thank you for correcting me. The destination is London, United Kingdom. Now, please provide me with the weight of the package. We consider the generated dialogue as Version 1 and generate three additional versions of the same dialogue, each resolving one of the feedback scenarios.
Resolving Feedback Scenarios
Figure 3 illustrates the idea. For each version, we first mask the affected system utterance and generate a replacement using the preceding dialogue context and task-specific information. Next, we drop the following two utterances, since they are directly related to the generation error. This way, the dialogue remains coherent and the conversation continues with the next regular user utterance [^7] We experimented with different ideas for resolving feedback scenarios (see Appendix C), but the naive approach described here turned out to be the most reliable..
Figure 3: Feedback dialogue generation. Each version solves one of the feedback scenarios from Version 1. See Appendix C (Figure 13) for an example dialogue.
We continue the process until all feedback scenarios have been resolved as in Version 4. For slot values, we only regenerate the annotations for the replaced system utterances in Version 2 to 4 and retain the other annotations from Version 1.
@@ 5 FEDI Analysis
FEDI consists of 8,852 dialogues, divided into 1,988 feedback-free dialogues, including 326 for testing, and 6,864 feedback dialogues (1,716 in four versions). The test dialogues were collected human-human by eight computer science students. In the following, we focus on the completeness of generated slot and intent annotations, the distribution of user emotions and the feedback scenarios represented in the dialogues. We provide additional statistical analysis in Appendix E, including split sizes and the distribution of demographic information. Details on recruitment, salary, procedure, and our experiences and findings from collecting and annotating dialogue data with humans vs. LLMs can be found in Appendix D.
@@
Table 9 shows the ratio of dialogues that provide all annotations for intent and slot values.
@@
Feedback
Gen. Test V1 V2 V3 V4
@@
Table 2: Ratio of dialogues that contain all annotations for related intent and slot values [^9] Hallucinated slot values, i.e., slot annotations with a value that does not occur in the respective utterance, were small in number and are not considered in the results.. For feedback-free dialogues, we distinguish generated (Gen.) and test dialogues (Test). The feedback dialogues are divided into versions, i.e. Version 1 (v1) to Version 4 (V4).
We observe a difference between Gen. and Test in the feedback-free dialogues, as the slot values often depend on the background stories. For example, with parcel shipping, if the user already has a shipping box, details about available shipping boxes are negligible. Human annotators consider this and omit slots if they are not required Zang et al. (2020). GPT-3.5 strictly follows our instructions, which include all slots as part of the task description. Question answering is less affected by this due to the more trivial annotation scheme (see Appendix A). In the feedback dialogues, the generated corrections sometimes do not contain all the required slot values. This is expected, because these dialogues focus on learning how to handle errors and feedback situations. We provide more findings as part of our manual analysis in Section 6.
@@
As expected, negative emotions are more common in feedback dialogues. Happiness in feedback dialogues is mostly observed when the system addresses the implicit user feedback. This is similar for Curiosity, although we also observe this emotion when the system suddenly changes the topic. While this emotion fits most dialogue context, it can also be the result of insufficient information in the emotion annotation instruction, as we only use the dialogue context as additional information and no further examples (see Appendix C).
@@
Figure 5 shows the distribution of user feedback types in relation to generation error types represented in the feedback scenarios of the feedback dialogues.
Figure 5: Distribution of user feedback types in relation to generation error types in feedback scenarios.
It shows that our approach for generating feedback scenarios mostly resulted in meaningful combinations of generation error and user feedback types. For example, Factually Incorrect is mostly addressed by Make Aware with Correction. Unclear Intention and Attribute Error are frequently addressed by Ask for Clarification and Repeat or Rephrase. The latter one is also frequently observed in combination with Ignore Question and Ignore Expectation errors, although Ignore and Continue is the most frequent user feedback to these generation error types.
6 Quality Control for FEDI
We asked two participants from our test data collection to assess and curate the intent, slot and emotion annotations in 480 feedback-free dialogues and the generation error and implicit user feedback type annotations in 380 feedback dialogues (see Appendix D.3 for the procedure). The dialogues were randomly sampled from the train and dev splits of FEDI. We used INCEpTION Klie et al. (2018) as a platform for this study. We calculate the inter-annotator agreement (IAA) using Krippendorff’s Alpha Krippendorff (2006) with a nominal weighting function (as provided in the platform). Table 3 shows the results [^11] Overall, 26 dialogues were reported as off-topic (13/480 feedback-free and 13/380 feedback). They are not considered in these results. The curated dialogues were not considered in our experiments but are included as a separate set in the published dataset..
Annotation Type Missing Changed IAA
Feedback-Free Dialogues Intent 0.06 0.35 0.90
Slot Values 0.56 0.19 0.83
User Emotions 0.02 0.81 0.91
Feedback Dialogues Generation Error Type 0.16 0.36 0.97
User Feedback Type 0.16 0.34 0.89
Overall, the ratio of dialogues with at least one missing annotation is rather low, except for slot annotations. We found that most of them are parcel shipping dialogues, which has a comparatively complex annotation scheme (see Appendix A). A detailed analysis revealed that an average of 1.8 annotations were added to these dialogues. For the dialogues with at least one changed annotation, annotators reported that in many of these cases placeholders, e.g., the slot name put in brackets ([shipping_box_name]), were used instead of the slot values from the dialogues. We attribute this to our observation from Section 5 (GPT-3.5 strictly follows the slot annotation scheme, even if the values are not in the dialogue). Emotions, whose perception is very subjective, are the most frequently changed annotation type (on average 2.09 times per affected dialogue), whereby the originally annotated emotion was sometimes not part of our taxonomy (hallucination). We provide further analysis of the impact of human curation on data quality in Appendix E.1.
Experiment Task Completion Quality Generation Accuracy
Inform Success Intent Acc. Slot Acc. Toxicity F1 BLEU BertScore
Flan-T5 Feedback-Free Flan-T5 86.7 85.9 54.8 60.9 52.7 0.02 45.0 20.0 88.3
   +User Emotions 83.9 83.2 61.2 58.3 57.5 0.02 46.7 21.0 88.9
   +Demographic Info. 87.0 86.0 33.5 29.3 54.5 0.03 43.2 18.4 87.7
   +User Emotions
   +Demographic Info.
85.3 85.1 43.9 36.7 56.4 0.02 44.2 19.1 88.1
Feedback    +Generation Error 96.8 92.7 72.5 76.7 56.9 0.02 41.4 19.8 87.8
   +User Feedback 96.6 94.1 69.0 76.2 56.3 0.02 41.3 19.3 87.6